A man walk from a point A and walk a distance of 8.0 km due northeast to point B from B he then walks 4.0 km due southeast to point x. Calculate the shortest distance between And x.


1
Expert's answer
2021-06-20T18:00:58-0400

Gives

AB distance move north east direction


Man walk from Ato B=8km


Bx distance moves sound east


Next man walk B to x =4km

"AX^2=BA^2+BX^2"

"AX=\sqrt{8^2+4^2}=\sqrt{80}=8.94m"